OpenStudy (anonymous):

In the school marching band the ratio of trumpet players to drummers was 5 to 2. If there were 6 drummers in the marching band, how many trumpet players were there?

OpenStudy (anonymous):

is there anything else to the question?

OpenStudy (anonymous):

@mathman123 i fixed the question

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

15 I think beautiful :)

OpenStudy (texaschic101):

5/2 = x/6 cross multiply 2x = 30 x = 15 I agree with mathman

OpenStudy (anonymous):

lol

OpenStudy (texaschic101):

think of these as equivalent fractions 5/2 * 3/3 = 15/6

OpenStudy (anonymous):

thank you both @mathman123 @texaschic101

OpenStudy (anonymous):

youre welcome hun